OpenStudy (anonymous):

x + y/y=4 and x+z/z=6, what is the value of y/z? A) 3/5 B) 2/3 C) 1 D) 3/2 E) 5/3

OpenStudy (anonymous):

any explaination :o?

OpenStudy (across):

It's simple algebra. We have: \[(x+y)/y=4\]and\[(x+z)/z=6\]Now solve for both y and z: \[x/4+y/4=y\]and\[x/6+z/6=z\]Then:\[x/4=(3/4)y\]and\[x/6=(5/6)z\]Isolate:\[x/3=y\]and\[x/5=z\]Finally, divide:\[(x/3)/(x/5)=5x/3x\]Check on my algebra; I did this in notepad...

OpenStudy (anonymous):

Lala do u know this? To graph the inequality y > 5x - 4, you would draw a solid line. A. True B. False

OpenStudy (anonymous):

B

OpenStudy (across):

95, try asking someplace else.

OpenStudy (across):

By the way, the x's cancel out.

OpenStudy (anonymous):

lalaly help me!

OpenStudy (anonymous):

ohh ty :) ^.^